OpenStudy (anonymous):

(*HELP PLEASE*) What is the solution to the system of equations represented by the two equations? y= (-1/2x) =4 y = 3/2x (***Both Equations are Fractions***) --- I know i'm supposed to set them equal to find the answer but I don't know how. Please walk me through it!

OpenStudy (campbell_st):

well can you check the 1st equation, it can't be what you have typed. is it \[y = \frac{-1}{2} x + 4\]

OpenStudy (anonymous):

Yes ^ Both equations are fractions.

OpenStudy (campbell_st):

because each are equal to y... the can be equated so you need to solve \[\frac{3}{2} x = -\frac{1}{2} x + 4\] add 1/2 x to both side of the equation then solve for x

OpenStudy (texaschic101):

if you do not like to mess with fractions, multiply the entire equation by 2...then solve for x